if 85% of a number is 97000, what is the number?
Answer by RAY100(1637)

Let n=number
.
85% of n = 97,000,,,,, per cent means per centum,,,or per hundred,,85% = 85/100=.85
.
.85 * n =97,000
.
n= 97,000/ .85 = 114,118
.
check
.
.85 * 114,118 = 97,000,,,ok